OpenStudy (anonymous):

Angie was working on solving the exponential equation 23^x = 6; however she isn't sure where to start. Describe to Angie how to solve this equation and how solving it would be different if the bases were equal. Please Help Explain This To Me! SUPER IMPORTANT!

OpenStudy (anonymous):

@abb0t @dumbcow @eliassaab @Hero @flvsguy

OpenStudy (dumbcow):

take log of both sides \[\ln (23^x) = \ln 6\] then using log identity .... ln a^n = n*ln a \[x \ln 23 = \ln 6\] \[x = \frac{\ln 6}{\ln 23}\]

OpenStudy (anonymous):

are the bases equal in this equation? Would they just be 10?

OpenStudy (dumbcow):

no bases are different .... 23 and 6 if bases are equal, then you can just set exponents equal to each other without using logs \[4^x = 16 = 4^2\] \[x = 2\]

OpenStudy (anonymous):

okay because this question has 2 parts: solve 23^x=6 with equal bases and then solve 23^x=6 with different bases.

OpenStudy (anonymous):

would those just be the answers just not in full depth? @dumbcow

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

@quietus

OpenStudy (anonymous):

@dumbcow HELP!